# An Erdős problem on random subset sums in finite abelian groups

### v1

@@ Abstract.
This work is an outcome of human–AI collaboration: the original qualitative proof was generated autonomously by ChatGPT-5.2 Pro, while the quantitative refinement was developed by the authors.
@@
random subset sums, Erdős problems, abelian groups, second moment method
@@ 1. Introduction
In 1973, Erdős [6, p. 127] conjectured that the $O(\log\log N)$ error term in (1) cannot be sharpened to $o(\log\log N)$. This conjecture is listed as Problem #543 on Bloom’s Erdős Problems website [4] in the following self-contained form.
@@
In this paper we confirm Erdős’ conjecture by answering Problem 1.1 in the negative. We restrict attention to prime orders $N=p$ (so $G$ has to be isomorphic to the cyclic group $\mathbb{F}_{p}$ of order $p$). This suffices to rule out the proposed bound $f(N)\leq\log_{2}N+o(\log\log N)$ for general $N$. The notion $\mathbb{F}_{p}$ usually stands for the field of $p$ elements, and here we only view it as an additive group. For any subset $A\subseteq\mathbb{F}_{p}$ we write
@@
1.1. Declaration of AI usage
The interest of this work lies not only in the mathematical results presented but also in the process by which it was obtained, with large language models serving as a major research assistant. For transparency and reproducibility, we provide the AI-assisted workflow below. [^1] For a public record of the development, see the second author’s comments posted at 13:12 on 21 Jan 2026 and 11:37 on 23 Jan 2026 in the forum thread at https://www.erdosproblems.com/forum/thread/543.
  1. (i)
    In an initial chat, we asked ChatGPT-5.2 Pro to search for a counterexample to the Erdős problem stated in [4] (i.e., the inequality in Problem 1.1). The model returned a proposed counterexample together with an accompanying argument.
  2. (ii)
    In a separate chat, we asked ChatGPT-5.2 Pro to verify the proof from the first session. It reported a serious gap that invalidated the argument as written.
  3. (iii)
    We then asked the second chat session to repair the argument. It produced a patched version that indeed disproves the inequality in Problem 1.1. The AI-generated proof, with minimal edits for readability, is available at [1].
  4. (iv)
    Motivated by the upper bound due to Erdős and Rényi [8], we observed that for prime $p$ one can obtain a quantitative lower bound, rather than only the qualitative disproof produced by the AI. By strengthening several steps of the AI-generated argument, we derived the quantitative lower bound for $f(p)$ stated in Theorem 1.2.
We emphasize that step (iv), as well as the writing of the paper, was carried out by the human authors. Related papers on AI-assisted solutions to Erdős problems have appeared recently; see, for example, Sothanaphan [13], Barreto et al. [3], and Feng et al. [9]. In our view, this work serves as an illustrative case study of AI-assisted use of the probabilistic method and linear algebra in combinatorics.
@@ 1.2. Paper organization
The rest of the paper is organized as follows. Section 2 collects notation and several auxiliary lemmas. In Section 3 we prove Theorem 1.2 by first establishing the key Poisson-like estimate, Proposition 3.1, and then combining it with a second-moment argument. Section 4 contains several concluding remarks.
@@ 2.1. Notation
Throughout, $\log$ denotes the natural logarithm and $\log_{2}x:=\log x/\log 2$. For a positive integer $k$, we define $k]:=\{1,\dots,k\}$. For $Y\in\mathbb{R}$ and $r\in\mathbb{N}$, let $(Y)_{r}:=Y(Y-1)\cdots(Y-r+1)$ denote the falling factorial.
@@ 2.2. Auxiliary lemmas
We now present several auxiliary lemmas that will be used repeatedly in the proof of the main theorem. We begin with two linear-algebraic observations for $\{0,1\}$-matrices. Let $\mathbb{Q}$ denote the field of rational numbers.
@@
Fix a constant $\beta$ with $0<\beta<\tfrac{1}{2}$. Let $p$ be prime and let $V\in\{0,1\}^{r\times k}$ with $r\leq(\log p)^{\beta}$. Then for all sufficiently large primes $p$,
@@
Let $V\in\{0,1\}^{r\times k}$ have distinct and nonzero rows. Let $W:=\mathrm{Col}(V)\subseteq\mathbb{Q}^{r}$ denote the column space of $V$ over $\mathbb{Q}$. Then its orthogonal complement $W^{\perp}$ contains no nonzero vector supported on at most two coordinates.
@@
Assume for contradiction that there exists some $y\in W^{\perp}\backslash\{0\}$ whose support has size at most $2$. We distinguish between the following two cases.
Case 1: $\mathrm{supp}(y)=\{i\}$. Then $y=ce_{i}$ for some $c\in\mathbb{Q}^{\times}$. Since $y\in W^{\perp}$, we have $\langle y,x\rangle=0$ for all $x\in W$, i.e. $c\,x_{i}=0$ for all $x\in W$, hence $x_{i}=0$ for all $x\in W$. In particular, taking $x$ to be any column of $V$, we obtain that the $i$th entry of every column is 0, so the $i$th row of $V$ is identically zero, contradicting the hypothesis.
Case 2: $\mathrm{supp}(y)=\{i,j\}$ with $i\neq j$. Then $y=\alpha e_{i}+\beta e_{j}$ for some $\alpha,\beta\in\mathbb{Q}^{\times}$. Again $y\in W^{\perp}$ implies
$\alpha x_{i}+\beta x_{j}=\langle y,x\rangle=0\qquad\text{for all }x\in W.$
Applying this to each column $v^{(t)}\in\mathbb{Q}^{r}$ of $V$ gives, for every $1\leq t\leq k$,
$\alpha\,V_{i,t}+\beta\,V_{j,t}=0.$
Equivalently, as vectors in $\mathbb{Q}^{k}$,
$\alpha\,V_{i,\ast}+\beta\,V_{j,\ast}=0,$
so $V_{i,\ast}=c\,V_{j,\ast}$ with $c:=-\beta/\alpha\in\mathbb{Q}^{\times}$, i.e. the $i$th and $j$th rows are proportional. Since $V_{j,\ast}\in\{0,1\}^{k}$ is nonzero, there exists an index $t$ with $V_{j,t}=1$, hence $c=V_{i,t}\in\{0,1\}$. As $c\neq 0$, we must have $c=1$, and therefore $V_{i,\ast}=V_{j,\ast}$, contradicting the assumption that the rows are distinct.
Hence no such $y$ exists in both cases, completing the proof of this lemma. ∎
@@
Choose a coordinate set $J\subseteq[r]$ with $|J|=d$ such that the projection $\pi_{J}:W\to\mathbb{Q}^{J}$ is an isomorphism. Thus each $x\in W$ is uniquely determined by $u:=x|_{J}\in\mathbb{Q}^{d}$, and for each $i\notin J$ there exists a homogeneous linear form $L_{i}:\mathbb{Q}^{d}\to\mathbb{Q}$ with $x_{i}=L_{i}(u)$. Hence
$W\cap\{0,1\}^{r}\ \cong\ \Bigl\{u\in\{0,1\}^{d}:\ L_{i}(u)\in\{0,1\}\ \forall i\in[r]\setminus J\Bigr\}.$
We claim that there exists some $i_{0}\in[r]\setminus J$ such that $L_{i_{0}}$ has at least two nonzero coefficients. Indeed, suppose for contradiction that each $L_{i}$ depends on at most one coordinate (equivalently, its coefficient vector has support size at most $1$). Since each $L_{i}$ is a homogeneous linear form on $\mathbb{Q}^{d}$, for every $i\in[r]\setminus J$ we then have either
$L_{i}(u)\equiv 0,$
or else there exist $j(i)\in J$ and $c_{i}\in\mathbb{Q}^{\times}$ such that
$L_{i}(u)=c_{i}\,u_{j(i)}\qquad\text{for all }u\in\mathbb{Q}^{d}.$
In the first case, $x_{i}=L_{i}(x|_{J})\equiv 0$ for all $x\in W$, hence $e_{i}\in W^{\perp}$ (a nonzero vector of support $1$). In the second case, for all $x\in W$ we have $x_{i}=L_{i}(x|_{J})=c_{i}x_{j(i)}$, so $x_{i}-c_{i}x_{j(i)}\equiv 0$ on $W$ and therefore $e_{i}-c_{i}e_{j(i)}\in W^{\perp}$ (a nonzero vector of support $2$). In either case we obtain a nonzero vector in $W^{\perp}$ supported on at most two coordinates, contradicting the hypothesis. Thus there must exist $i_{0}\in[r]\setminus J$ for which $L_{i_{0}}$ has at least two nonzero coefficients.
Fix such $i_{0}$ and write $L_{i_{0}}(u)=\sum_{\ell=1}^{d}\alpha_{\ell}u_{\ell}$ with at least two $\alpha_{\ell}\neq 0$. Let $Z$ be uniform on $\{0,1\}^{d}$. Conditioning on all coordinates except two indices $s\neq t$ with $\alpha_{s},\alpha_{t}\neq 0$, the random variable $L_{i_{0}}(Z)$ becomes $C+aX+bY$ where $X,Y$ are independent $\mathrm{Bernoulli}(1/2)$ and $a,b\neq 0$. Thus $C+aX+bY$ takes four equiprobable values; in the degenerate case $a=\pm b$ it takes three values with masses $1/4,1/2,1/4$. In either case,
@@
$k=\Bigl\lfloor\log_{2}p+c\log\log p\Bigr\rfloor,\qquad M:=2^{k}-1,\qquad\mbox{ and }\qquad\lambda:=\frac{M}{p},$
@@
Write $\theta_{p}:=\log_{2}p+c\log\log p-k\in[0,1)$. Then
$2^{k}=2^{\log_{2}p+c\log\log p-\theta_{p}}=p\,2^{c\log\log p}\,2^{-\theta_{p}}=p\,(\log p)^{c\log 2}\,2^{-\theta_{p}}=p\,(\log p)^{\alpha}\,2^{-\theta_{p}}.$
Thus $2^{k}/p$ lies between $(1/2)(\log p)^{\alpha}$ and $(\log p)^{\alpha}$. Since $M=2^{k}-1\sim 2^{k}$, we obtain $\lambda=M/p=\Theta((\log p)^{\alpha})$. The remaining claims follow immediately. ∎
@@
Write
$(mM)_{r}=\prod_{j=0}^{r-1}(mM-j)=(mM)^{r}\prod_{j=0}^{r-1}\Bigl(1-\frac{j}{mM}\Bigr).$
Since $0\leq j\leq r-1<(\log p)^{\beta}$, we have $j/(mM)=o(1)$ uniformly in $j$, and thus the fact $\log(1-x)=-x+O(x^{2})$ leads to
$\sum_{j=0}^{r-1}\log\Bigl(1-\frac{j}{mM}\Bigr)=-\frac{1}{mM}\sum_{j=0}^{r-1}j+O\Bigl(\frac{1}{(mM)^{2}}\sum_{j=0}^{r-1}j^{2}\Bigr).$
Using $\sum_{j<r}j=O(r^{2})$ and $\sum_{j<r}j^{2}=O(r^{3})$, we obtain
$\sum_{j=0}^{r-1}\log\Bigl(1-\frac{j}{mM}\Bigr)=O\Bigl(\frac{r^{2}}{M}\Bigr)+O\Bigl(\frac{r^{3}}{M^{2}}\Bigr)=o(1),$
because $r\leq(\log p)^{\beta}$ while $M=2^{k}-1\geq p^{1+o(1)}\to\infty$. Therefore
$(mM)_{r}=(mM)^{r}\exp(o(1))=(mM)^{r}\,(1+o(1)).$ (2)
Dividing by $p^{r}$ yields
$\frac{(mM)_{r}}{p^{r}}=m^{r}\Bigl(\frac{M}{p}\Bigr)^{r}(1+o(1)).$
By Lemma 2.4, we have $M/p=\lambda=\Theta((\log p)^{\alpha})$ and thus $\log(M/p)=\Theta(\log\log p)$. So
$\log\mathopen{}\mathclose{{\left(\Bigl(\frac{M}{p}\Bigr)^{r}}}\right)=r\log(M/p)=O((\log p)^{\beta}\log\log p)=o(\log p),$
implying that $(M/p)^{r}=p^{o(1)}$. Combining the above estimates gives
$\frac{(mM)_{r}}{p^{r}}=m^{r}\,p^{o(1)},$
as desired. ∎
@@ 3. Proof of Main Result
In this section we prove Theorem 1.2 in three stages. First, in Section 3.1, we reduce the problem to an independent (i.i.d.) model. The proof of the theorem itself (Section 3.2) is based on a second-moment analysis, which relies on a crucial Poisson-like estimate (Proposition 3.1). The proof of this proposition, which constitutes our main technical contribution, is given in Section 3.3 and proceeds by applying Bonferroni inequalities as well as estimating factorial moments of an associated Poisson-like random variable.
@@ 3.1. Reducing to an independent model
In the independent (i.i.d.) model, let $a_{1},\dots,a_{k}$ be i.i.d. uniform on $\mathbb{F}_{p}$ (allowing repetitions) and set
$A_{\mathrm{iid}}:=\{a_{1},\dots,a_{k}\}\subseteq\mathbb{F}_{p},\qquad\mathbb{P}_{\mathrm{iid}}(F):=\mathbb{P}\bigl(F(A_{\mathrm{iid}})\bigr).$
@@
Conditioned on $E$, the unordered set $A_{\mathrm{iid}}$ is uniformly distributed over $\binom{\mathbb{F}_{p}}{k}$, so $A_{\mathrm{iid}}\mid E$ has the same distribution as $A_{\mathrm{subset}}$. By the law of total probability, for any event $F$ we have
@@
Applying (4) to the event $F=\{\Sigma(A)=\mathbb{F}_{p}\}$ yields
@@
$\mathbb{P}\bigl(X_{B}=0\bigr)=e^{-m\lambda}\,(1+o(1))\qquad(p\to\infty),$
@@
Assuming Proposition 3.1, we now complete the proof of Theorem 1.2.
@@
By Proposition 3.1 with $m=1$, for each fixed $x\in\mathbb{F}_{p}^{\times}$ we have
$\mathbb{P}(X_{x}=0)=e^{-\lambda}(1+o(1)).$
@@
$\mathbb{E}U=\sum_{x\in\mathbb{F}_{p}^{\times}}\mathbb{P}(X_{x}=0)=(p-1)e^{-\lambda}(1+o(1))=pe^{-\lambda}(1+o(1)).$ (8)
Similarly, for distinct $x\neq y$ in $\mathbb{F}_{p}^{\times}$, Proposition 3.1 (applied with $m=2$ and $B=\{x,y\}$) yields
$\mathbb{P}(X_{x}=0,\ X_{y}=0)=e^{-2\lambda}(1+o(1)),$
and hence we can derive that
$\mathbb{E}[U(U-1)]=\sum_{\begin{subarray}{c}x,y\in\mathbb{F}_{p}^{\times}\\ x\neq y\end{subarray}}\mathbb{P}(X_{x}=0,X_{y}=0)=(p-1)(p-2)e^{-2\lambda}(1+o(1))=p^{2}e^{-2\lambda}(1+o(1)).$ (9)
From (8)–(9) we obtain
$\frac{\mathbb{E}[U(U-1)]}{(\mathbb{E}U)^{2}}=\frac{p^{2}e^{-2\lambda}(1+o(1))}{p^{2}e^{-2\lambda}(1+o(1))}=1+o(1),$
and thus
$\frac{\mathrm{Var}(U)}{(\mathbb{E}U)^{2}}=\frac{\mathbb{E}U+\mathbb{E}[U(U-1)]-(\mathbb{E}U)^{2}}{(\mathbb{E}U)^{2}}=\frac{1}{\mathbb{E}U}+\mathopen{}\mathclose{{\left(\frac{\mathbb{E}[U(U-1)]}{(\mathbb{E}U)^{2}}-1}}\right)=\frac{1}{\mathbb{E}U}+o(1).$
By Lemma 2.4 we have $\lambda=o(\log p)$, so $\log(\mathbb{E}U)=\log p-\lambda+o(1)\to\infty$ and hence $\mathbb{E}U\to\infty$. Therefore $\mathrm{Var}(U)/(\mathbb{E}U)^{2}\to 0$, and Chebyshev’s inequality gives
$\mathbb{P}(U=0)\leq\mathbb{P}\bigl(|U-\mathbb{E}U|\geq\mathbb{E}U\bigr)\leq\frac{\mathrm{Var}(U)}{(\mathbb{E}U)^{2}}=o(1).$
Combining this with (7) and (5), we conclude
@@
which proves Theorem 1.2. ∎
@@
Fix $m\in\{1,2\}$ and $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m$. We estimate the factorial moments of $X_{B}$ by comparing them with those of a Poisson random variable of mean $m\lambda$, and then apply these bounds in the following Bonferroni-type inequalities (which arise from the inclusion-exclusion principle on factorial moments; see, e.g., [2, Lemma 2.1] and [5, Theorem 8.2.1(b)]). For every integer-valued random variable $Z\geq 0$, it holds that

### v2

@@ Abstract.
random subset sums, abelian groups, Erdős problems.
@@ 1. Introduction
In 1973, Erdős [4, p. 127] conjectured that the $O(\log\log N)$ error term in (1) cannot be sharpened to $o(\log\log N)$. This conjecture is listed as Problem #543 on Bloom’s Erdős Problems website [2] in the following self-contained form.
@@
In this paper we confirm Erdős’ conjecture by answering Problem 1.1 in the negative. We restrict attention to prime orders $N=p$ (so $G$ has to be isomorphic to the cyclic group $\mathbb{F}_{p}$ of order $p$). This suffices to rule out the proposed bound $f(N)\leq\log_{2}N+o(\log\log N)$ for general $N$. The notation $\mathbb{F}_{p}$ usually stands for the field of $p$ elements, and here we only view it as an additive group. For any subset $A\subseteq\mathbb{F}_{p}$ we write
@@
1.1. Proof overview and comparison with earlier work
The Erdős–Rényi upper bound [6] sets the benchmark for the problem. We now discuss the lower-bound side and the main ideas of the present proof.
A closely related earlier work is that of Erdős and Hall [5], who worked in an i.i.d. model on a finite abelian group $G$ of order $n$. For every $r\geq 0$, they defined $d(r)$ to be the number of group elements having exactly $r$ representations as subset sums of $k$ independent random elements. In particular, $d(0)$ denotes the number of missed elements (i.e., elements not representable as subset sums of these random elements). Write $\mu=2^{k}/n$. They proved a general theorem showing that, under the assumption that the number of elements of each fixed order in $G$ is $o(n)$, if $k=\log_{2}n+O(1)$, then for every fixed integer $r\geq 0$,
$d(r)=\bigl(1+o(1)\bigr)\,ne^{-\mu}\frac{\mu^{r}}{r!}$
with probability tending to $1$ as $n\to\infty$. Their proof is moment-theoretic, based on the analysis of fixed-order moments and variances of the representation counts, combined with character-sum estimates.
For the present problem, however, the decisive quantity is the single statistic $d(0)$. Erdős and Hall [5] also proved an additional result for cyclic groups: there exists an absolute constant $b>0$ such that $d(0)>0$ with probability tending to $1$ whenever $\mu<b\log\log n$. Since $\mu=2^{k}/n$, this is equivalent to
$k<\log_{2}n+\log_{2}(b\log\log n)=\log_{2}n+O(\log\log\log n).$
Their proof again relies on the moment method, but uses the additional structure of cyclic groups to extend the fixed-moment analysis to a growing range of moment orders. For broader background on sets of multiples and probabilistic group theory, see Hall [7, Chapter 4].
The present paper is close in spirit to [5], but technically quite different. We work only in the prime cyclic case $G=\mathbb{F}_{p}$, and we do not seek asymptotics for all fixed multiplicities $d(r)$. Instead, after reducing the random $k$-subset model to an i.i.d. model, we focus on the number $d(0)$ of missed elements and reduce the problem to estimating the probability $\mathbb{P}_{B}$ that all elements of a prescribed subset $B\subseteq\mathbb{F}_{p}\setminus\{0\}$ with $|B|=1$ or $2$ are missed. More precisely, if $X_{B}$ denotes the number of indexed subset sums landing in $B$, then we require a Poisson-type estimate for $X_{B}$ with mean $|B|\lambda$, where
$\lambda=\frac{2^{k}-1}{p}\to\infty.$
In our application we take $k=\lfloor\log_{2}p+c\log\log p\rfloor$, which implies
$\lambda\asymp(\log p)^{\alpha}\qquad\alpha=c\log 2\in(0,1/2).$
We then estimate the desired probability $\mathbb{P}_{B}:=\mathbb{P}(X_{B}=0)$ using Bonferroni-type inequalities, which require uniform control of the factorial moments of $X_{B}$ up to a truncated level $R\asymp(\log p)^{O(1)}\to\infty$; see the proof of Proposition 3.1. This approach is technically different from that of Erdős–Hall [5].
Another new technical ingredient is a quantitative bound for low-rank incidence matrices (see Lemma 3.4), obtained via a Boolean-cube intersection bound. Together, these ingredients allow us to raise the earlier cyclic obstruction from the $\log\log\log p$-scale to the $\log\log p$-scale.
@@ 1.2. Paper organization
The rest of the paper is organized as follows. Section 2 collects notation and several auxiliary lemmas. In Section 3 we prove Theorem 1.2 and Corollary 1.3 by first establishing the key Poisson-like estimate, Proposition 3.1, and then combining it with a second-moment argument. Section 4 contains several concluding remarks.
@@ 2.1. Notation
Throughout, $\log$ denotes the natural logarithm and $\log_{2}x:=\log x/\log 2$. For a positive integer $k$, we define $k]:=\{1,\dots,k\}$. For $Y\in\mathbb{R}$ and $r\in\mathbb{N}$, let $(Y)_{r}:=Y(Y-1)\cdots(Y-r+1)$ denote the falling factorial, with the convention $(Y)_{0}:=1$.
@@ 2.2. Auxiliary lemmas
We now present several auxiliary lemmas that will be used repeatedly in the proof of the main theorem. We begin with two linear-algebraic observations for $\{0,1\}$-matrices.
Let $\mathbb{Q}$ denote the field of rational numbers. We equip $\mathbb{Q}^{r}$ with the standard inner product, and write $W^{\perp}$ for the orthogonal complement of a subspace $W\subseteq\mathbb{Q}^{r}$.
@@
Fix a constant $\beta$ with $0<\beta<\tfrac{1}{2}$. Let $r,k\geq 1$, and let $V\in\{0,1\}^{r\times k}$ with $r\leq(\log p)^{\beta}$. Then for all sufficiently large primes $p$,
@@
Let $r,k\geq 1$, let $V\in\{0,1\}^{r\times k}$ have distinct and nonzero rows, and let $W$ denote the column space of $V$ over $\mathbb{Q}$. Then $W^{\perp}$ contains no nonzero vector supported on at most two coordinates.
@@
Let $0\neq y\in W^{\perp}$, and write $I=\operatorname{supp}(y)\subseteq[r]$. Then for every $x\in W$,
$\sum_{i\in I}y_{i}x_{i}=0.$
In particular, applying this to each column $v^{(t)}\in W$ of $V$, we get
$\sum_{i\in I}y_{i}V_{i,t}=0\qquad(1\leq t\leq k).$
Hence $\sum_{i\in I}y_{i}\,R_{i}=0$, where $R_{i}\in\mathbb{Q}^{k}$ denotes the $i$th row of $V$. Thus the rows of $V$ indexed by $I$ are linearly dependent over $\mathbb{Q}$.
If $|I|=1$, this says that one row of $V$ is zero, contradicting the hypothesis. If $|I|=2$, then two rows of $V$ are proportional. Since these rows are nonzero vectors in $\{0,1\}^{k}$, proportionality forces them to be equal, again contradicting the hypothesis that the rows are distinct. ∎
@@
Choose a coordinate set $J\subseteq[r]$ with $|J|=d$ such that the projection $\pi_{J}:W\to\mathbb{Q}^{J}$ is an isomorphism. Thus each $x\in W$ is uniquely determined by $u:=x|_{J}\in\mathbb{Q}^{d}$, and for each $i\notin J$ there exists a homogeneous linear form $L_{i}:\mathbb{Q}^{d}\to\mathbb{Q}$ with $x_{i}=L_{i}(u)$. For $j\in J$, also set $L_{j}(u)=u_{j}$. Hence
$|W\cap\{0,1\}^{r}|=\mathopen{}\mathclose{{\left|\mathopen{}\mathclose{{\left\{u\in\{0,1\}^{d}:\ L_{i}(u)\in\{0,1\}\ \forall i\in[r]\setminus J}}\right\}}}\right|.$
The assumption on $W^{\perp}$ implies that the forms $L_{1},\dots,L_{r}$ are pairwise linearly independent. Indeed, if $L_{i}\equiv 0$, then $x_{i}=0$ for all $x\in W$, so $e_{i}\in W^{\perp}$, where $e_{i}$ denotes the $i$th standard basis vector of $\mathbb{Q}^{r}$; and if $L_{i}=cL_{j}$ for some distinct $i,j\in[r]$ and some $c\in\mathbb{Q}^{\times}$, then $x_{i}-cx_{j}=0$ for all $x\in W$, so $e_{i}-ce_{j}\in W^{\perp}$. In particular, for every $i\in[r]\setminus J$, the form $L_{i}$ has at least two nonzero coefficients.
Fix any $i_{0}\in[r]\setminus J$ and write $L_{i_{0}}(u)=\sum_{\ell=1}^{d}\alpha_{\ell}u_{\ell}$ with at least two $\alpha_{\ell}\neq 0$. Let $Z$ be uniform on $\{0,1\}^{d}$. Conditioning on all coordinates except two indices $s\neq t$ with $\alpha_{s},\alpha_{t}\neq 0$, the random variable $L_{i_{0}}(Z)$ becomes $C+aX+bY$ where $X,Y$ are independent $\mathrm{Bernoulli}(1/2)$ and $a,b\neq 0$. Thus $C+aX+bY$ takes four equiprobable values; in the degenerate case $a=\pm b$ it takes three values with masses $1/4,1/2,1/4$. In either case,
@@
$k:=\Bigl\lfloor\log_{2}p+c\log\log p\Bigr\rfloor,\qquad M:=2^{k}-1,\qquad\mbox{ and }\qquad\lambda:=\frac{M}{p},$ (2)
@@
By (2), we have
$2^{k}\asymp p(\log p)^{c\log 2}=p(\log p)^{\alpha}.$
Since $M=2^{k}-1\sim 2^{k}$, we get $\lambda=M/p=\Theta((\log p)^{\alpha})$. As $\alpha<1$, this gives $\lambda=o(\log p)$, and hence $e^{\lambda}=p^{o(1)}$. ∎
@@
By Lemma 2.4,
$M\asymp p(\log p)^{\alpha}.$
Since $r\leq(\log p)^{\beta}$ with $\beta<1/2$, we have $r^{2}/M=o(1)$. Therefore
$(mM)_{r}=(mM)^{r}\prod_{j=0}^{r-1}\Bigl(1-\frac{j}{mM}\Bigr)=(mM)^{r}\exp\Bigl(O\Bigl(\frac{r^{2}}{M}\Bigr)\Bigr)=(mM)^{r}(1+o(1)),$ (3)
uniformly for $m\in\{1,2\}$. Dividing by $p^{r}$, we get
$\frac{(mM)_{r}}{p^{r}}=m^{r}\Bigl(\frac{M}{p}\Bigr)^{r}(1+o(1))=m^{r}\lambda^{r}(1+o(1)).$
Finally, by Lemma 2.4,
$\log(\lambda^{r})=r\log\lambda=O((\log p)^{\beta}\log\log p)=o(\log p),$
so $\lambda^{r}=p^{o(1)}$. Hence
$\frac{(mM)_{r}}{p^{r}}=m^{r}\,p^{o(1)}.\qed$
Remark 2.6.
The hypotheses in Lemmas 2.1 and 2.5 can be weakened. We have chosen not to optimize them here, since the stated forms already suffice for all subsequent applications in this paper.
@@ 3. Proof of Main Result
In this section we prove Theorem 1.2 and Corollary 1.3 in three stages. First, in Section 3.1, we reduce the problem to an independent (i.i.d.) model. The proof of the theorem itself (Section 3.2) is based on a second-moment analysis, which relies on a crucial Poisson-like estimate (Proposition 3.1). The proof of this proposition, which constitutes our main technical contribution, is given in Section 3.3 and proceeds by applying Bonferroni inequalities as well as estimating factorial moments of an associated Poisson-like random variable.
@@ 3.1. Reducing to an independent model
In the independent (i.i.d.) model, let $a_{1},\dots,a_{k}$ be i.i.d. uniform on $\mathbb{F}_{p}$, and set
$A_{\mathrm{iid}}:=(a_{1},\dots,a_{k}).$
Thus $A_{\mathrm{iid}}$ is an ordered $k$-tuple; when convenient, we also regard it as a multiset, so repetitions are allowed. For the i.i.d. sample $A_{\mathrm{iid}}=(a_{1},\dots,a_{k})$, we define
$\Sigma(A_{\mathrm{iid}}):=\Bigl\{\sum_{i\in I}a_{i}:\ I\subseteq[k]\Bigr\}.$
This depends only on the underlying multiset of $a_{1},\dots,a_{k}$. For any event $F$ depending only on the underlying multiset $A_{\mathrm{iid}}$, we write
$\mathbb{P}_{\mathrm{iid}}(F):=\mathbb{P}\bigl(F(A_{\mathrm{iid}})\bigr).$
@@
Conditioned on $E$, the multiset $A_{\mathrm{iid}}$ is uniformly distributed over $\binom{\mathbb{F}_{p}}{k}$, so $A_{\mathrm{iid}}\mid E$ has the same distribution as $A_{\mathrm{subset}}$. Let $F$ be any event in the i.i.d. model that depends only on the multiset $A_{\mathrm{iid}}=\{a_{1},\dots,a_{k}\}$. Then the corresponding event in the subset model satisfies
$\mathbb{P}_{\mathrm{subset}}(F)=\mathbb{P}_{\mathrm{iid}}(F\mid E).$
By the law of total probability, for any such event $F$ we have
@@
Applying (5) to the event $F=\{\Sigma(A_{\mathrm{iid}})=\mathbb{F}_{p}\}$ yields
@@
$\mathbb{P}\bigl(X_{B}=0\bigr)=(1+o(1))e^{-m\lambda}\qquad(p\to\infty),$
@@
Assuming Proposition 3.1, we now complete the proofs of Theorem 1.2 and Corollary 1.3.
@@
In this proof, we use a standard second-moment argument; see, for example, [1]. By Proposition 3.1 with $m=1$, uniformly for all $x\in\mathbb{F}_{p}^{\times}$,
$\mathbb{P}(X_{x}=0)=(1+o(1))e^{-\lambda}.$
@@
$\mathbb{E}U=\sum_{x\in\mathbb{F}_{p}^{\times}}\mathbb{P}(X_{x}=0)=(1+o(1))(p-1)e^{-\lambda}.$
Similarly, for distinct $x\neq y$ in $\mathbb{F}_{p}^{\times}$, Proposition 3.1 (applied with $m=2$ and $B=\{x,y\}$) yields, uniformly in such pairs,
$\mathbb{P}(X_{x}=0,\ X_{y}=0)=(1+o(1))e^{-2\lambda},$
and hence
$\mathbb{E}U^{2}$ $=\sum_{x\in\mathbb{F}_{p}^{\times}}\mathbb{P}(X_{x}=0)+\sum_{\begin{subarray}{c}x,y\in\mathbb{F}_{p}^{\times}\\ x\neq y\end{subarray}}\mathbb{P}(X_{x}=0,\ X_{y}=0)$
$=(1+o(1))(p-1)e^{-\lambda}+(1+o(1))(p-1)(p-2)e^{-2\lambda}.$
Hence
$\mathrm{Var}(U)=\mathbb{E}U^{2}-(\mathbb{E}U)^{2}=O(pe^{-\lambda})+o(p^{2}e^{-2\lambda}).$
By Lemma 2.4, $\lambda=o(\log p)$, so
$\mathbb{E}U\asymp pe^{-\lambda}\to\infty.$
Therefore
$\mathrm{Var}(U)=O(\mathbb{E}U)+o\bigl((\mathbb{E}U)^{2}\bigr)=o\bigl((\mathbb{E}U)^{2}\bigr).$
By Chebyshev’s inequality,
$\mathbb{P}(U=0)\leq\frac{\mathrm{Var}(U)}{(\mathbb{E}U)^{2}}=o(1).$
Combining this with (8) and (6), we obtain
@@
as required. ∎
Proof of Corollary 1.3, assuming Proposition 3.1.
Let $\varepsilon>0$ be an arbitrary constant, and set $c=\frac{1}{2\log 2}-\varepsilon$. By Theorem 1.2, if $k=\lfloor\log_{2}p+c\log\log p\rfloor$, then
$\mathbb{P}_{\mathrm{subset}}\bigl(\Sigma(A)=\mathbb{F}_{p}\bigr)\to 0\qquad(p\to\infty).$
In particular, for all sufficiently large primes $p$, this probability is less than $\tfrac{1}{2}$. By the definition of $f(p)$, it follows that $f(p)>k$. Therefore
$f(p)\geq k+1\geq\log_{2}p+\mathopen{}\mathclose{{\left(\frac{1}{2\log 2}-\varepsilon}}\right)\log\log p,$
and the result follows. ∎
@@
Fix $m\in\{1,2\}$ and $B\subseteq\mathbb{F}_{p}^{\times}$ with $|B|=m$. We estimate the factorial moments of $X_{B}$ by comparing them with those of a Poisson random variable of mean $m\lambda$, and then apply these bounds in the following Bonferroni-type inequalities (which arise from the inclusion-exclusion principle on factorial moments; see, e.g., [8, Eq. (1.276)] and [3, Exercises 1.6.9–1.6.10]). For every nonnegative integer-valued random variable $Z$ with finite support, it holds that